OpenStudy (one098):

What is the inverse function of f(x)=2x+9

OpenStudy (freckles):

You need to solve y=2x+9 for x as a first step.

OpenStudy (igreen):

\(f(x) = 2x + 9 \rightarrow f^{-1}(x) = \dfrac{x-9}{2}\)

OpenStudy (igreen):

@One098

OpenStudy (one098):

Okay, thanks

OpenStudy (anonymous):

f(x) = 2x + 9 let y = 2x + 9

OpenStudy (anonymous):

interchange x and y x = 2y + 9

OpenStudy (one098):

Ok thanks!

OpenStudy (anonymous):

now make y the subject of the formula

OpenStudy (anonymous):

x = 2y + 9 x - 9 = 2y (x - 9)/2 = y

OpenStudy (anonymous):

f‾¹(x) = (x - 9)/2 and u get this :)